OpenStudy (anonymous):

Help ; 3x + 4x - x = x + 6

OpenStudy (anonymous):

7x-x=x+6 6x=x+6 6x-x=6 5x=6 x=6/5

OpenStudy (anonymous):

Help ; 16 - (2y - 4) = 4y

OpenStudy (anonymous):

16 - 2y + 4 = 4y -2y - 4y = -20 -6y = -20 y = 10/3 or\[3\frac{ 1 }{ 3 }\]